OpenStudy (anonymous):

Can anyone please help me to solve this? y/y+1-1/y=1

OpenStudy (anonymous):

\[\frac{y}{y+1}-\frac{1}{y}=1 \]\[\frac{y}{y+1}-\frac{1}{y}-1=0\]Combine the fractions in the LHS.\[\frac{-2 y-1}{y (y+1)}=0 \]\[y=-\frac{1}{2} \]

OpenStudy (anonymous):

how did you get 2y?

OpenStudy (campbell_st):

easiest way is to multiply each term by y(y+1) this will eliminate the fractions then after cancelling common factors y^2 - (y +1) = y(y+1) simplifying -2y - 1 = 0 -2y = 1 y = -1/2

OpenStudy (anonymous):

Thanks, both of you